OpenStudy (anonymous):

Find the derivative of each function defined as follows. y=12^3-8x^2+7x+5

OpenStudy (anonymous):

\[y=12x^3-8x^2+7x+5\]

jimthompson5910 (jim_thompson5910):

hint: \[\Large y = x^n\] \[\Large y^{\prime} = nx^{n-1}\]

jimthompson5910 (jim_thompson5910):

So for example \[\Large y = x^7\] \[\Large y^{\prime} = 7x^6\]

OpenStudy (anonymous):

i dont understand

jimthompson5910 (jim_thompson5910):

Another example \[\Large y = 3x^8\] \[\Large y^{\prime} = 8*3x^{8-1}\] \[\Large y^{\prime} = 24x^{7}\]

OpenStudy (anonymous):

can you help me step by step with this question

jimthompson5910 (jim_thompson5910):

This is the power rule. It says "to derive x^n, you pull down the exponent n and multiply it with the coefficient and then subtract the exponent" Check out this page http://www.mathscoop.com/calculus/derivatives/the-power-rule.php

OpenStudy (anonymous):

So the first one will be 36x^2

jimthompson5910 (jim_thompson5910):

that is correct what about the next one?

OpenStudy (anonymous):

-16x

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

how about the next one

OpenStudy (anonymous):

7x

OpenStudy (anonymous):

or 7

jimthompson5910 (jim_thompson5910):

7x = 7x^1 so the derivative of 7x^1 is 1*7x^0 = 7*1 = 7

jimthompson5910 (jim_thompson5910):

So put all together, the derivative of \[\Large y=12x^3-8x^2+7x+5\] is \[\Large y=36x^2-16x+7\]

OpenStudy (anonymous):

what happens to the 5

jimthompson5910 (jim_thompson5910):

the derivative of any constant is 0

jimthompson5910 (jim_thompson5910):

so it effectively goes away

OpenStudy (anonymous):

okay now whats next

jimthompson5910 (jim_thompson5910):

that's it, the final answer is posted above

OpenStudy (anonymous):

oh okay i get it. Thank you very much. Ill contact you if i get stuck with a problem.

OpenStudy (anonymous):

how do i turn this into a derivative x/12

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\Large y = \frac{x}{12}\] \[\Large \frac{d}{dx}[y] = \frac{d}{dx}\left[\frac{x}{12}\right]\] \[\Large \frac{d}{dx}[y] = \frac{d}{dx}\left[\frac{1}{12}*x\right]\] \[\Large \frac{d}{dx}[y] = \frac{1}{12}\frac{d}{dx}\left[x\right]\] \[\Large \frac{d}{dx}[y] = \frac{1}{12}\frac{d}{dx}\left[x^{1}\right]\] \[\Large \frac{d}{dx}[y] = \frac{1}{12}*1*x^{1-1}\] \[\Large \frac{d}{dx}[y] = \frac{1}{12}*1*x^{0}\] \[\Large \frac{d}{dx}[y] = \frac{1}{12}*1*1\] \[\Large \frac{d}{dx}[y] = \frac{1}{12}\] \[\Large y^{\prime} = \frac{1}{12}\]

OpenStudy (anonymous):

\[y=8\sqrt{x}+6x ^{3/4}\]

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

I need help with this one

jimthompson5910 (jim_thompson5910):

rewrite \[\Large \sqrt{x}\] as \[\Large x^{1/2}\]

jimthompson5910 (jim_thompson5910):

then use the power rule to derive

jimthompson5910 (jim_thompson5910):

\[\Large y=8\sqrt{x}+6x ^{3/4}\] \[\Large \frac{d}{dx}[y]=\frac{d}{dx}[8\sqrt{x}+6x ^{3/4}]\] \[\Large y^{\prime}=\frac{d}{dx}[8x^{1/2}]+\frac{d}{dx}[6x^{3/4}]\] \[\Large y^{\prime}=8\frac{d}{dx}[x^{1/2}]+6\frac{d}{dx}[x^{3/4}]\] \[\Large y^{\prime}=8*\frac{1}{2}*x^{1/2-1}+6*\frac{3}{4}*x^{3/4-1}\] \[\Large y^{\prime}=4*x^{-1/2}+\frac{9}{2}*x^{-1/4}\] \[\Large y^{\prime}=\frac{4}{x^{1/2}}+\frac{9}{2x^{1/4}}\] \[\Large y^{\prime}=\frac{4}{\sqrt{x}}+\frac{9}{2\sqrt[4]{x}}\] You can optionally rationalize the denominator, but that's not necessary in my opinion.